# Emerging Quadrature Lattices of Kerr Combs

### v1

@@ 4.2 The finite model and analysis of the robustness of non-detuned quadrature lattice modes
where the matrices $U$ and $V$ are unitary and $D\left(\omega\right)=\text{diag}\left\{d_{1}\left(\omega\right),...,d_{N}\left(\omega\right)|d_{1}^{-1}\left(\omega\right),...,d_{N}^{-1}\left(\omega\right)\right\}$ corresponds to squeezing and anti-squeezing values of the different quadrature supermodes. Next we aim to show more broadly the impact of group velocity dispersion and asymmetry in dispersion relative to the comb on the multi-mode states. In Fig. 7(a-b) we present a typical complex eigenvalue band structure for a quadrature lattice with $D_{2}\neq 0$, setting $D_{2}/2\pi=1.2~{}MHz$, $\mathcal{R}=0.225$, and $Q_{i}=1.74\cdot 10^{6}$, which are experimental parameters also used in Fig. 2 of the main text. Clearly, dispersion compresses the spectrum bandwidth of the quadrature supermode and the number of participating supermodes. By focusing on the most relevant supermodes, we observe that APT symmetry persists: Supermodes are either purely real (un-detuned) or purely imaginary (detuned), with the purely real ones exhibiting higher gain.
@@
Figure 7: Existence of non-detuned state in the presence of non-zero $D_{2}$ and $D^{\prime}_{1}$ a. Eigenvalue solution of the quadrature Hamiltonian $\mathcal{M}$ for a 2-FSR Kerr comb with side band drop ratio of $\mathcal{R}=0.225$, and $D_{2}/2\pi=1.2~{}MHz$. b. Zoom-in on the quadrature supermodes with higher-gain showing that all of them are undetuned (imaginary value is 0.) c-d. Same as a. and b. only with $D^{\prime}_{1}/2\pi=2.8~{}MHz$. e.-f. Numerically counting the un-detuned states as a function of underlying GVD (D2) and detuning of the comb for the symmetric case of $D^{\prime}_{1}=0$ and the non-symmetric case $D^{\prime}_{1}/2\pi=2.8~{}MHz$
@@
Introducing a mismatch between the cavity FSR and the Kerr-comb frequency spacing eliminates lattice symmetry, yielding more general complex eigenvalue solutions for the $\mathcal{M}$ matrix. When the mismatch is sufficiently small, as depicted in Fig. 7(c-d), we observe that some supermodes remain with purely real eigenvalues while others do not. Figure 7(e,f) shows the existence of un-detuned (imaginary eigenvalue is 0) supermodes as a function of $D_{2}$ and constant pump detuning. Figure 7(e) enumerates supermodes with purely real eigenvalues for $D^{\prime}_{1}=0$, while Fig. 7(f) counts the supermodes when $D^{\prime}_{1}/2\pi=2.8~{}MHz$.
4.3 Additional data on fabricated device and experimental setup
Our devices are etched from a 550 nm thick 4H-SiC-on-Insulator [43] and capped with SiO2. Microheaters are patterned via liftoff of 100 nm platinum with a 5 nm titanium adhesion layer. Coupling to and from the chip is achieved using lensed fibers, with collection efficiency of up to 50%.
Figure 8: Micro-ring resonator: fabrication and dispersion. a. Optical image of the Kerr ring resonator alongside the smaller on-chip filter ring and heaters. b. Integrated dispersion of the TE mode of the Kerr ring resonator (blue) and fit to $D_{2}=1.2~{}MHz,D_{1}=152.7$ GHz (red). Dispersion defects are avoided mode crossings with other mode families.
The ring resonators exhibit intrinsic quality factors of up to 8 million, with typical values around 1.5 million. The resonators exhibit a $D_{2}/2\pi$ of 1.2MHz, and an FSR of 153 GHz. Additionally, the ring is multimode, causing mode splitting through modal interactions (Fig. 8(b), which allows a broad range of narrow-bandwidth states to be observed in the devices, such as 2-FSR primary combs (Fig. 9(a-b)), as well as the transition to 1-FSR combs and soliton crystals. The heaters modify the mode-splitting properties, thereby altering the detuning at threshold and consequently the Kerr comb’s frequency spacing $\Delta\Omega$. Figure 9 presents the experimental evolution of the comb for the symmetric case. The comb starts as a 2-FSR primary comb (Fig. 9(a-b)) and undergoes an undetuned (single peak) threshold to a 1-FSR phase locked comb (Fig. 9(c)). Following the threshold, a second detuned (double peak) threshold occurs followed by additonal thresholds and an rf beatnote in the comb (Fig. 9(d,e)). Finally, the comb abruptly turns to a 2-FSR soliton crystal (Fig. 9(f)).
Figure 9: Kerr comb for different pump detuning - a single evolution. a.-b. A 2-FSR primary comb, where b. is closer to the secondary threshold (the pump is at a longer wavelength) than a. c. Tuning the pump to longer wavelengths, the 2-FSR comb evolves to a single phase locked 1-FSR comb corresponding with Fig. 5(c-d). d.-e. Additional thresholds creating additional 1-FSR combs with subcombs corresponding with Fig. 5(d). f. Red tuning the pump further results in a 2-FSR soliton crystal (corresponding with the measurements in Fig. 4)
Next, we provide a more detailed description of the experimental setup (Fig. 10(a-b)). We utilize two tunable CW telecom lasers named CW laser 1 (Velocity, 1520-1570 nm, TLB-6728) and CW laser 2 (Toptica CTL1500: 1460-1570nm). The lasers are split into two channels. The first channel (upper channel in Fig. 10(b)) receives light from laser 1 and after amplification through an EDFA excites the Kerr resonator. The Kerr resonator outputs two channels: the first carries the quantum signal with the filtered Kerr comb directly to the homodyne unit, while the second channel contains the Kerr comb passing thorough the filter ring to a photodiode.
The second channel (bottom in Fig. 10(b)), receives light either from CW laser 1 or CW laser 2 or both. CW laser 1 is used for measuring modes -1 and 1, while CW laser 2 allows measuring other modes but is not phase-locked to CW laser 1.
4.4 Multimode Homodyne detection
To measure the maximal squeezing and anti-squeezing of the multimode below-threshold light, balanced homodyne detection must be performed with a multi-tone local oscillator (LO). The intensity and relative phase of each line of the LO should match the maximally-squeezed supermode of the squeezed vacuum [10]. In this work, the multi-tone LO is generated from a 20 GHz EOM comb [48] driven by the same pump laser. The low bandwidth of a single phase modulator (or a few cascaded phase modulators) requires either working with relatively large micro-resonators or only partial overlap with the maximally squeezed supermode. Large micro-resonators pose additional challenges, including higher operation powers (for the same intrinsic quality factor) and a lower $D_{2}$ for the same resonator cross-section. The theory and experimental validation we present in this work does not overlap optimally with the squeezed supermode, but is sufficient to capture the frequency-dependence of the most squeezed states. Therefore we are able to use a two-tone local oscillator despite the fact that the connectivity expands beyond two modes. We found that the two-tone local oscillator measures significantly stronger quadrature variance signal than a single-tone local oscillator by improving the overlap with the squeezed mode. Additionally, the two-tone symmetric LO may measure a time-independent phase relative to the signal in the absence of active phase locking of the EOM comb to the repetition rate of the Kerr comb [49].
We now provide further mathematical details on the multi-tone homodyne measurement.
To calculate the noise variance spectrum for a multi-tone local oscillator at the repetition rate of the comb [33], we can define a supermode decay operator
$\ \hat{L}(t)=\sum_{j}\alpha_{j}\hat{a}_{j}(t)$ (11)
whose Hermitian conjugate creates a photon in a superposition of azimuthal modes $\hat{a}_{j}(t)$ across the frequency comb. $\alpha_{j}=|\alpha_{j}|e^{i\phi_{j}}$ defines the amplitude and phase composition of the supermode. Its associated supermode output operator is
$\ \hat{L}_{\text{out}}(t)=\sum_{j}\alpha_{j}\hat{b}_{\text{out},j}(t)$ (12)
where $\hat{b}_{\text{out},j}(t)$ is the output operator of mode $\hat{a}_{j}(t)$. The corresponding Hermitian quadrature operators of $\hat{L}_{\text{out}}$ are:
$\ \hat{L}_{\text{out}}^{(+)}(t)=\frac{1}{\sqrt{2}}\big{(}\hat{L}_{\text{out}}^{\dagger}(t)+\hat{L}_{\text{out}}(t)\big{)}$ (13)
$\ \hat{L}_{\text{out}}^{(-)}(t)=\frac{i}{\sqrt{2}}\big{(}\hat{L}_{\text{out}}^{\dagger}(t)-\hat{L}_{\text{out}}(t)\big{)}$ (14)
The noise variance is then defined as
@@ 4.4 Multimode Homodyne detection
The maximally-squeezed supermodes can be obtained from the columns of $U(\omega)$ (Eq. 10), which defines the linear combination of quadratures $(q_{1}(t),...,q_{n}(t)|p_{1}(t),...,p_{n}(t))^{T}$ for a given supermode. When $U(\omega)$ is real, then the composition for the $k$th maximally-(anti)squeezed supermode at $\omega$ may be written:
$\ |\alpha_{j}|=\sqrt{U_{j,k}(\omega)^{2}+U_{j+n,k}(\omega)^{2}}$ (16)
$\ \phi_{j}=\operatorname{atan2}{(U_{j+n,k}(\omega),U_{j,k}(\omega))}$ (17)
The quantum (anti)-squeezing is therefore significant when the local oscillator composition overlaps significantly with a multimode state that has large value $d_{i}\left(\omega\right)$. In this approach maximum quadrature variance can be obtained through this method for the undetuned squeezing. Since the undetuned squeezing occurs at $\omega=0$, the value of its eigenvector is real $U\left(0\right)\in\mathbb{R}$ and a time-independent local oscillator (in the frame rotating at the repetition rate of the comb) will measure maximal (anti)-squeezing. In the case of detuned squeezing, $U\left(\omega\right)$ is complex and a stationary local oscillator will not necessarily measure maximal squeezing.
However, when measuring a symmetric double peak configuration one may obtain maximal squeezing [49]: when a local oscillator is equidistant between a nondegenerate squeezed state, measurement at the sideband frequency probes both modes. If the double peak is symmetric, we are summing two peaks with opposite Fourier components. Of course, as shown in Methods 4.6 the double peak quadrature variance spectrum can be asymmetric.
4.5 Additional data on LLE simulation
In order to support the theoretical conclusions and experimental observations, we preformed LLE simulations of a 2-FSR Kerr comb. The LLE simulates the classical comb dynamics [44], and allows us to focus on specific aspects, some not previously addressed in the experiment. Particularly, the squeezing structures, both detuned and non-detuned, significantly influence the formation of secondary combs. Hence, studying the primary comb to secondary comb transitions in a classical framework is valuable in the context of the findings presented in this work.
We start by simulating a perturbed dispersion yielding a comb similar to that of our experiment. The dispersion perturbation in our setup changes in response to heating. For the same reason, the perturbation also changes as a function of power in the cavity (i.e it is changing during the evolution of the comb). However, since the effects we describe do not require this behavior, we can observe the main features under constant dispersion. We use experimentally extracted parameters of $D_{2}/2\pi=1.2MHz$ and extract the perturbation parameters from the measured dispersion. For simplicity, we only use the perturbed modes near the pump modes $-2,-1,0,1,2$. For quality factors, we use a constant value of 1.83 million, which we reduce at modes distant 17 FSR away from the pump to avoid pair generation there. The normalized pump power is $f=5.93$ where $f^{2}=8g_{0}\kappa_{c}P_{in}/\kappa^{3}\hbar\omega_{0}$ and $P_{in}$ is the input power, $\kappa_{c}$ is the outcoupling power loss, and $\omega_{0}$ is the pump angular frequency.
Importantly, the formation of our 2-FSR comb is not a direct result of the anomalous dispersion, but it is formed locally due to the dispersion defect (see Methods  4.3). This means that the initial comb formation may be governed by adjusting the dispersion of sites $-2,0,2$. The detuning value of the threshold to a 1-FSR comb depends on the dispersion of the odd modes and the perturbation of modes $-1,1$. By adjusting these parameters, we are able to simulate a comb evolution that is qualitatively similar to what we observe in the experiment.
Figures 5(a-b) and 11(a) show the comb evolution in the simulation, which is similar to our experiment (Fig. 9). The process begins with a 2-FSR primary comb which transitions to a 1-FSR secondary comb. Figure 11 demonstrates the transitions from one rolling Turing pattern to another. Since this transition is un-detuned, the first 1-FSR comb is a single phase locked comb and does not have sub combs.
Figure 11: Intracavity field in real space for different regimes in the comb formation - LLE with perturbed dispersion - undetuned threshold. a. Transmission as a function of continuous pump detuning (toward longer wavelengths). This plot is the same as in Fig. 5(a) but the shaded areas mark different regimes plotted in (b-d). b. Intracavity field in real space for the transition between the primary comb (2-FSR) and the secondary comb (1-FSR) corresponding to the shaded red region in a. c. Same as b but for the transition from the secondary comb (1-FSR) to a second detuned 1-FSR comb which oscillates (shaded green region). d. The region of a 2-FSR soliton crystal (shaded blue region).
The next transition involves the threshold crossing of a detuned 1-FSR comb, which produces an rf beatnote with the existing 1-FSR comb. The oscillations in this state are evident in both the transmission and the intra-cavity field (Fig. 11(c)). Finally the system transitions to a 2-FSR soliton crystal (Fig. 11(d)). To better understand the unique symmetric transition from a 2-FSR comb to a single phase locked 1-FSR comb, we explore the comb state at a pump detuning close to this threshold at 0.115 GHz.
Figure 12: Examination of the simulated threshold between an un-detuned 2-FSR comb to an un-detuned 1-FSR comb at 0.115 GHz detuning. a. Spectrum of the 2-FSR primary Kerr comb (red) and light in odd modes (green). b. Intracavity field in real space (rolling turing) c. $g^{(2)}$ correlations of the below threshold light. d. Eigen-spectrum of the comb showing that an un-detuned supermode crosses threshold, and that the spectrum corresponds to the partially symmetric case. e. The squeezing structure, showing the single peak which correspond with the un-detuned supermode that crosses threshold in d. f. The spectrum of the most squeezed state
Figure  12(a-b) shows the comb from Fig. 11 for the detuning of 0.115 GHz in both the frequency (mode) space and real physical space respectively. Using the obtained comb amplitudes and the hot-cavity dispersion, we calculate the $g^{(2)}$ correlations of each cavity-mode $\mu$ with each cavity-mode $\nu$, which resembles the experimental $g^{(2)}$ in Fig. 3(b). Next, we calculate the quadrature Hamiltonian $M$ associated with the 2-FSR comb at 0.115GHz. The eigenspectrum is plotted in Fig. 12(d), showing that the non-detuned state (purely real eigenvalue) is approaching threshold. Correspondingly, the squeezing structure is a single peak (Fig. 12(e)), localized near the pump (Fig. 12(f)).
In addition to the perturbed dispersion case, we wish to explore in simulation the case of un-perturbed dispersion with a $D_{2}>0$ term. Throughout this work, a perturbation in dispersion somewhat detached the general dispersion of the cavity from the frequency spacing of the 2-FSR Kerr comb, allowing for more possible behaviors.
Figure 13: Simulation of a 2-FSR primary comb transition to a 1-FSR secondary comb in a unperturbed dispersion. a. Transmission vs pump-detuning of the 2-FSR comb without dispersion perturbation. b. spectrum of the 2-FSR Kerr comb (red) in an unperturbed dispersion. c. Eigen-spectrum showing 40 supermodes. In this case all the modes are detuned (purely imaginary eigenvalues) while the Hamiltonian is symmetric. d. Zoom-in on the spectrum in c. e. The squeezing structure, showing the double peak which correspond with the detuned state that crosses threshold in c,d f.. The spectrum of the most squeezed state
However, when a comb is formed without perturbation, the general dispersion and the comb’s frequency spacing are more inherently linked, reducing the number of possible outcomes. While this case is conceptually simpler than the perturbed dispersion case, it can be more challenging to realize due to the high anomalous dispersion that is required for the 2-FSR comb to form (together with a relatively small FSR). Additionally, high anomalous dispersion causes the squeezed supermodes to localize in four modes, reducing the relevance of the lattice model description (Fig. 6). Nevertheless, this case does highlight some aspects which are relevant to the main case studied in this work. To generate a 2-FSR comb, we use high anomalous dispersion of $D_{2}/2\pi=86.4~{}MHz$, an intrinsic Q of 3.66 million and coupling Q of 4.25 million. These parameters result in a 2-FSR primary comb, which we pump at 12.55mW with CW light at 1545nm. This comb reaches threshold at a pump detuning of $-0.04~{}GHz$ (Fig. 13(a)). At this pump detuning value, even though the dispersion is symmetric and $D^{\prime}_{1}=0$, the threshold is still detuned (double peak). The symmetry implies that the eigenvalues are purely real or purely imaginary. Thus, if all the states are purely imaginary, there are no un-detuned states in existence, leading to a double peak threshold. Correspondingly, the 2-FSR comb in Fig. 13b, for the detuning marked by the orange dot in Fig. 13(a), has a fully detuned spectrum (Fig. 13(c,d)). Such a spectrum is associated with a detuned threshold, i.e, a squeezing spectrum that peaks away from zero. The localization that high $D_{2}$ dispersion causes can be seen in the spectrum of the supermode in Fig. 13(f). The localization occurs for all supermodes, as Bragg scattering is not efficient away from the center of the comb.
4.6 Single supermode dominance
As we demonstrate in the theoretical and experimental measurements of this work, the quantum noise in the different cavity modes is dominated by one or two supermodes. Additionally, the squeezing frequency spectrum is uniform for each supermode, meaning that the below threshold light, has similar anti-squeezing in all cavity modes in our experiment. In this section we show experimentally how a single supermode dominates the spectrum close to threshold. We focus on the non-symmetric case, which shows distinct two peaks, and perform a single-tone homodyne measurement in each cavity mode. The results are presented in Fig. 14, which shows how similar two peaks cross the threshold for each cavity mode.
Figure 14: Threshold crossing of detuned supermode by individually homodyning modes -5 to 7. a.-g. Extension of the data in Fig. 5(e,f), showing additional 5 homodyne measurements of the threshold crossing of a double peak (detuned) threshold, showing the dominance of a single supermode in all modes
Note that the two peaks appear both in the below threshold light (lower line plots in each sub-figure) and above threshold light (upper line-plots).
[…]
@@ 4.6 Single supermode dominance
Acknowledgements We gratefully acknowledge discussions with Edwin Ng, Ryotatsu Yanagimoto, and Kiyoul Yang. This work is funded by the Defense Advanced Research Projects Agency under the QuICC program and by the NSF QuSeC-TAQS Award ID 2326792, as well as the Vannevar Bush Faculty Fellowship from the US Department of Defense. E.L acknowledges the Yad Hanadiv Rothschild fellowship, and the Zuckerman institute Zuckerman fellowship. Part of this work was performed at the Stanford Nanofabrication Facility (SNF) and the Stanford Nano Shared Facilities (SNSF). S. F. acknowledge the support of a MURI project for the U. S. Air Force Office of Scientific Research (Grant No. FA9550-22-1-0339). We thank NGK Insulators, Ltd. for 4H-SiCOI substrates used to fabricate devices in this work.

### v3

@@ 4.6 The finite model analysis
where the matrices $U$ and $V$ are unitary and $D\left(\omega\right)=\text{diag}\left\{d_{1}\left(\omega\right),...,d_{N}\left(\omega\right)|d_{1}^{-1}\left(\omega\right),...,d_{N}^{-1}\left(\omega\right)\right\}$ corresponds to squeezing and anti-squeezing values of the different quadrature supermodes. The eigenvectors (columns of $U\left\{\omega\right\}$) are morphing with $\omega$ which is detected within the RF spectrum. Thus, the optimization of the local oscillator of the multimode homodyne changes for the same supermode as a function of $\omega$.
Next we would address the OPO thresholds and comb transitions in the context of the complex eigenvalues of the finite model. The eigenvectors of $\mathcal{M}$ are not orthogonal under the usual (Hermitian) inner product which is why the eigenvalues are complex. The complexity of the eigenvalue close but before threshold has profound implications on threshold behavior and secondary comb formation.
First, complex eigenvalues near threshold result in the amplification of oscillating supermodes up to the secondary Kerr comb threshold. The oscillations of the vacuum fluctuations become a double peak in the RF spectrum at each cavity mode leading to multiple RF beat notes after threshold. Thus, if a supermode near threshold has an eigenvalue that is purely real (for example: due to the APT symmetry) RF beat notes will not appear, which will halt the formation of beat notes and consequently chaos until another supermode crosses the OPO threshold. We show this process experimentally in the supplementary section 3. As a side note: APT symmetry does not guarantee that a supermode with purely real eigenvalue will cross threshold, since an entire spectrum (all the supermodes) can also be purely imaginary in the APT case. In this scenario, all supermodes will oscillate and will approach threshold with equal gain until one supermode will break this symmetry and cross threshold. This scenario is simulated and discussed in section 2 of the Supplementary Material.
Second, measuring the squeezing close to threshold when the modes are oscillating (and non-orthogonal) leads to sub-optimal squeezing measurement, even for optimized local oscillator [15].
Next we aim to show more broadly the impact of group velocity dispersion and asymmetry in dispersion relative to the comb on the multi-mode states. We calculate the quadrature eigenspectrum of the sub-threshold light generated by a 2-FSR Kerr comb with fixed $A_{\mu}$ profile ($\mathcal{R}=2.25$ as previously) and varying $D^{\prime}_{1}$. The confinement of the lattice comes from the introduction of non-zero group velocity dispersion $D_{2}$ and a mismatch between the comb repetition rate and the mismatch $\tilde{D}^{\prime}_{1}$. Note that $\tilde{D}^{\prime}_{1}$ is the unitless frequency difference (divided by $g_{0}A_{0}^{2}$) between the cavity FSR and the Kerr comb’s frequency spacing.
For $D_{2}$ and the rest of the parameters we use the value derived from our experiment: $D_{2}/2\pi=1.2~{}MHz$. As mentioned $\tilde{D}^{\prime}_{1}$ varies as a function of the microheater voltage; therefore, we studied three cases: $\tilde{D}^{\prime}_{1}=0,3,8$ which correspond to the fast, slow, and non-oscillating regimes observed in Fig. 4(e). Additionally, unlike in an infinite lattice, a threshold exists in a finite lattice. To account for this, we include a constant loss value $\kappa$ corresponding to an intrinsic quality factor of 1.74 million. We note that while this value does not alter the qualitative behavior of the dynamics, it does affect it quantitatively. Setting $D_{2}/2\pi=1.2~{}MHz$, $\mathcal{R}=2.25$, and $Q_{i}=1.74\cdot 10^{6}$, Fig. 11(a) shows the non-oscillating case in which $D^{\prime}_{1}=0$, while Fig. 11b(c) are oscillating with $D^{\prime}_{1}/2\pi=3(8)MHz$. We adjust the overall comb power in the plots in Fig. 11 to be close to threshold for detuning $\Delta{\omega}_{0}=2g\sum_{\mu}\left|A_{\mu}\right|^{2}$.
To compute the eigenspectrum for a finite model we use Eq. 1 in a finite quadrature basis $(p_{1}...p_{N},q_{1}...q_{N})$ where $p_{i}$ and $q_{i}$ are the quadratures of mode $i$, and $N$ is the number of participating cavity modes. We obtain the finite dynamical matrix $\mathcal{M}$ which is non-Hermitian. To study threshold behavior we introduce $\Gamma=\kappa/2I_{2N\times 2N}$, where $\kappa$ is the uniform intensity loss rate.
In Fig. 11 we analyze the eigenspectrum of $\mathcal{M}-\Gamma$ while varying the parameter $D^{\prime}_{1}$, taking into account all orders of $\alpha$. We observe three distinct regimes. For $D^{\prime}_{1}=0$ (Fig. 11(a)), the eigenvalues of the different supermodes are either purely real or have a real part that equals $-\kappa/2$ resulting from their initial purely real eigenvalue shifted uniformly by $\kappa/2$. Interestingly, due to the multimode nature of the system, even when the symmetry is removed (i.e. when $D^{\prime}_{1}\neq 0$) some states close to threshold remain with imaginary values of 0 (i.e the transition region, see Fig. 11(b)). For the highest $D^{\prime}_{1}$ in Fig. 11(c), the modes close to threshold all have a double peak structure (i.e., imaginary part is non zero).
Figure 11: Dispersion symmetry and the properties of supermode squeezed states. a.-c. The real (blue dots) and imaginary (red dots) parts of the eigenvalues of the amplified vacuum supermodes driven by a 2-FSR comb with different values of $D^{\prime}_{1}$: 0, 3 and 8 MHz respectively. For larger $D^{\prime}_{1}$ the purely real eigenvalues gradually become complex. Highlighted quadrature supermodes are those close to threshold. d.-g. The frequency($\omega$)-dependent squeezing spectrum of the supermodes close to threshold (highlighted). Each line represents the supermode spectrum for a linearly swept intra cavity power, and the supermodes can be seen to narrow as they approach threshold. The supermode in f. does not narrow asymptotically because e. reaches threshold first. h.,i. The supermode composition as a function of cavity modes of the multimode squeezed states, for the supermodes in e. and f. respectively (colours correspond to the frequencies).
Each eigenvalue in Fig. 11(a-c) represents a multimode squeezed state, residing on the odd-numbered modes. As depicted for example in Fig. 3(b) under the assumptions taken so far the squeezing spectrum of each supermode uniformly evolves around each cavity mode [14]. The spectral structure around each cavity mode is frequency dependent squeezing, a concept which recently started to be utilized in metrology application through noise spectrum engineering [58].
To compute the multimode frequency-dependent squeezing structure, we calculate the supermodes of $\mathcal{M}-\Gamma$ for Eq. 10, using the Bloch-Messiah decomposition. Specifically, the squeezing RF spectrum of the dominant supermode that reaches the threshold is given by $d_{i}^{-1}(\omega)$ where $i$ corresponds to the supermode that has the largest real part of the eigenvalue. The RF structure of the supermodes with the largest real part (gain) in Fig. 11(a-c) is presented in Fig. 11(d-g). Supermodes with the highest gain will dominate in total photon number over other supermodes and have the narrowest peaks, being closest to the OPO threshold. The OPO threshold occurs when $Re(\xi)=0$, where $\xi$ denotes the eigenvalue of the supermode with the highest real part. The frequency spacing in the double peak structure is determined by the imaginary value of the corresponding supermode by $\Delta=2Im{\xi}$. Therefore, supermodes with imaginary value zero, exhibit a single peak structure (“non-oscillating” with respect to the frame of reference of the Kerr comb), while supermodes with imaginary value that is not zero exhibit double-peak spectrum (“oscillating” with respect to the frame of reference of the Kerr comb).
The transition between the two cases shows how the transition between a dominant double peak supermode and single peak supermode is not continuous but is the result of the singe peak supermode increasing in parametric gain compared to the double peak. The transition therefore exhibits both double-peak and single peak evolution for different supermodes with high gain simultaneously. The different plotted lines in each spectral shape in Fig 11(d-g) represent the squeezing for different distances from threshold (by changing the overall power of the Kerr comb). Together with an RF frequency-dependent structure, each supermode has a spectral structure in the cavity mode basis, shown for the transition region in Fig. 11(h-i). The spectral structure plotted here is the absolute value of the column vector that is extracted from $V\left(\omega=0\right)$. The bandwidth of these supermodes is predominantly determined by the values of $D^{\prime}_{1}$ and $D_{2}$.
@@
Next, we wish to discuss the robustness of the non-oscillating supermode. In Fig. 12(a) we present a typical complex eigenvalue band structure for the same experimental parameters as before: $D_{2}/2\pi=1.2~{}MHz$, $\mathcal{R}=2.25$, and $Q_{i}=1.74\cdot 10^{6}$. Clearly, dispersion compresses the spectrum bandwidth of the quadrature supermode and the number of participating supermodes. By focusing on the most relevant supermodes, we observe that APT symmetry persists: Supermodes are either purely real (non-oscillating) or purely imaginary (oscillating), with the purely real ones exhibiting higher gain.
@@
Figure 12: Existence of non-detuned state in the presence of non-zero $D_{2}$ and $D^{\prime}_{1}$ a. Top: eigenvalue solution of the quadrature Hamiltonian $\mathcal{M}$ for a 2-FSR Kerr comb with side band drop ratio of $\mathcal{R}=2.25$, and $D_{2}/2\pi=1.2~{}MHz$. Bottom: Zoom-in on the quadrature supermodes with higher-gain showing that all of them are undetuned (imaginary value is 0.) b. Same as a. only with $D^{\prime}_{1}/2\pi=2.8~{}MHz$. c.,d. Numerically counting the un-detuned states as a function of underlying GVD (D2) and detuning of the comb for the symmetric case of $D^{\prime}_{1}=0$ and the non-symmetric case $D^{\prime}_{1}/2\pi=2.8~{}MHz$
@@
Introducing a mismatch between the cavity FSR and the Kerr-comb frequency spacing eliminates lattice symmetry, yielding more general complex eigenvalue solutions for the $\mathcal{M}$ matrix. When the mismatch is sufficiently small, as depicted in Fig. 12(b), we observe that some supermodes remain with purely real eigenvalues while others do not. Figure 12(c,d) shows the existence of non-oscillating (imaginary part of the eigenvalue is 0) supermodes as a function of $D_{2}$ and constant pump detuning. Figure 12(c) enumerates supermodes with purely real eigenvalues for $D^{\prime}_{1}=0$, while Fig. 12(d) counts the supermodes when $D^{\prime}_{1}/2\pi=2.8~{}MHz$.
@@
Acknowledgements We gratefully acknowledge discussions with Edwin Ng, Ryotatsu Yanagimoto, and Kiyoul Yang. This work is funded by the Defense Advanced Research Projects Agency under the QuICC program and by the NSF QuSeC-TAQS Award ID 2326792, as well as the Vannevar Bush Faculty Fellowship from the US Department of Defense and AFOSR Award number: FA9550-23-1-0248. E.L acknowledges the Yad Hanadiv Rothschild fellowship, and the Zuckerman institute Zuckerman fellowship. Part of this work was performed at the Stanford Nanofabrication Facility (SNF) and the Stanford Nano Shared Facilities (SNSF). S. F. acknowledge the support of a MURI project for the U. S. Air Force Office of Scientific Research (Grant No. FA9550-22-1-0339). We thank NGK Insulators, Ltd. for 4H-SiCOI substrates used to fabricate devices in this work.